# Prompting Medical Large Vision-Language Models to Diagnose Pathologies by Visual Question Answering

### v1

@@ Abstract
Keywords: Medical Visual Question Answering, Large Vision-Language Models, Prompt Engineering
@@ 1 Introduction
Research on Large Language Models (LLMs) has yielded astonishing results in recent years. LLMs with billions of parameters have achieved outstanding abilities in a wide range of application scenarios (OpenAI, 2022; OpenAI et al., 2023; Chiang et al., 2023). The success of LLMs has quickly extended into the Vision-Language (VL) domain. Large Vision-Language Models (LVLMs) are built upon LLMs by training adapters that project visual features into features that can be interpreted by LLMs (Li et al., 2023b; Zhang et al., 2023; Liu et al., 2023b). Visual Question Answering (VQA) is an essential skill of LVLMs, and VQA accuracy serves as a test metric for most of these models (Li et al., 2023b; Zhang et al., 2023; Zhu et al., 2023; Liu et al., 2023b). LVLMs have been pretrained on medical datasets (Li et al., 2023a; Liu et al., 2023c; Singhal et al., 2023) and they have been tested on medical VQA tasks (Lau et al., 2018; He et al., 2020). These Medical LVLMs (MLVLMs) have been able to answer questions regarding the imaging modalities, organs, and abnormalities depicted by the input medical scans.
Unfortunately, however, “hallucination” has been a major problem for LVLMs. This refers to the generation of content that is contradictory to the input images. Hallucination can be measured via VQA. One may ask the model questions regarding the existence of objects in the input image(s) and the hallucination level is assessed as the percentage of correctly answered questions. VQA can also potentially serve for medical image diagnosis. Users pose questions regarding a pathology and the MLVLM responds based on its analysis of the medical scans. However, most of the available datasets involve simple questions such as “what is the modality of this image” and “what is the organ/tissue in this image”. MLVLMs have yet to be thoroughly evaluated on VQA accuracy across a broad spectrum of pathologies. Additionally, general VQA models are usually tested by the commonly known accuracy: the percentage of correctly answered questions, which is an unsuitable measure for medical VQA. Medical image classification metrics such as the Precision, Recall, and F1 are more suitable for the evaluation of medical VQA models. Several strategies have been explored to enhance the question answering of LLMs/LVLMs, including chain-of-thought prompting (Zheng et al., 2023), self-consistency (Wang et al., 2023), and retrieval-based augmentation (Caffagni et al., 2024). All these methods involve fine-tuning the models, which is expensive. Training-free methods to improve the VQA accuracy of MLVLMs are desirable.
@@
    We introduce a low-cost weak learner model as a reference for LLaVA-Med, and this effectively reduces the false positive (FP) answers.
@@ 2 Related Work
LVLMs are built upon LLMs. A pretrained visual encoder extracts the visual features and an adapter module projects the extracted features to ones that can be understood by the LLM. Models of this type include those by Liu et al. (2023b), Zhu et al. (2023), and Zhang et al. (2023). During training, the visual encoder and the LLM are usually fixed. VQA is an essential skill of LVLMs. Given an input image, the models should be able to answer questions correctly regarding that image.
@@
Strategies to mitigate hallucination in LVLMs mainly fall into two categories: prompt engineering and model improvement. Regarding the former, Liu et al. (2023a) leverage visual instructions constructed from the bounding box information in the input image to prompt LLMs. Zheng et al. (2023) use a chain of thought scheme to prompt the models to perform step-by-step visual-language reasoning like humans, which eventually leads to the correct answers. Wang et al. (2023) generate multiple chains of thought and use the one with the majority vote as the answer. Caffagni et al. (2024) prompt the model with explanations of the terms in questions. With regard to the model improvement strategy for reducing hallucination, Sun et al. (2023b) improve the visual and text feature alignment through reinforcement learning. Leng et al. (2023) propose a contrastive decoding strategy to reduce reliance on pretrained knowledge. Favero et al. (2024) and Zhao et al. (2024) also focus on the inference stage and propose specialized decoding strategies to mitigate hallucination. Other strategies for reducing hallucination have been proposed. For example, Zhou et al. (2024) design a post-processing model to detect hallucinated objects and rephrase the generated answers, and Sun et al. (2023a) adapt a reinforcement learning strategy that uses human evaluation of the hallucination level to improve the model.
@@
There are two approaches to assessing hallucination in LVLMs. The first is VQA. The ground truth information of the input images is leveraged to construct questions regarding the existence of objects in the images (e.g., “Is there a black cat in the image?”), as well as questions about objects which do not exist in the images. The models are evaluated in terms of the percentage of correctly answered questions. Metrics of this type include POPE (Li et al., 2023c), CIEM (Hu et al., 2023), and NOPE (Lovenia et al., 2023). The second approach is to use pre-designed prompts from which the models produce various generations that are then evaluated. Examples include CHAIR (Rohrbach et al., 2018), which counts the hallucinated objects in generated image captions, and MMHAL-BENCH (Sun et al., 2023b), which uses GPT-4 (OpenAI et al., 2023) to compare the generations with human answers and determine the propensity toward hallucination.
@@
For MLVLMs, given a medical scan, models such as LLaVA-Med (Li et al., 2023a) and Med-PALM (Liu et al., 2023c) are able to answer questions regarding the types of modalities, the scanned organs, and medical indicators such as opacity. They have demonstrated good performance on medical VQA datasets such as VQA-RAD (Lau et al., 2018), SLAKE (Liu et al., 2021), and Path-VQA (He et al., 2020). However, most medical questions in existing datasets are simple. MLVLMs have not been tested on a broader range of complex pathologies.
@@ 3.1 Prompting With Detailed Explanations
Prompt templates for a number of pathologies are listed in Section A.
Figure 3: Example of prompting an MLVLM for medical VQA using both pathology explanations and reference predictions from a weak learner.
@@ 3.2 Prompting With Detailed Explanations and Weak Learners
One can nevertheless enable MLVLMs to benefit by leveraging small models trained on re-sampled datasets. Our method resembles multiagent LLM systems, such as Du et al. (2023), where multiple LLMs debate each other and hallucination can be corrected by referring to the generated outputs of other models. Given that traditional image classifiers are smaller, it is feasible to train multiple small classifiers each of which is trained on re-sampled datasets of a particular pathology. Those models can be further fine-tuned to optimize a single aspect, such as fewer false positives (FP) or fewer false negatives (FN). The classifiers are applied to the medical images and return preliminary predictions. These predictions are selectively included in the prompts as references for the MLVLM. Hence, MLVLMs can benefit indirectly from the nuanced understanding that these specialized models can provide. This method is meaningful because clinicians usually must balance the trade-off between overtreatment and undertreatment when making healthcare decisions. For instance, they may prefer models having a low FP rate if the cost of overtreatment is higher than that of undertreatment.
@@ 4.1 Datasets
For our main testing regimen, we selected the five pathologies in the Chexpert Competition (Irvin et al., 2019), Atelectasis, Cardiomegaly, Consolidation, Edema, and Pleural Effusion.
@@ 4.2 Implementation Details
Table 3: The Prompt Templates (PTs). {target} is the pathology cited in the questions. {explanation} contains a pathology explanation among those listed in Section A. {n} is the probability associated with the weak learner.
@@
As will be justified by our experiments, our weak learner is designed to suppress FP predictions. To this end, we use the pretrained ResNet50 (He et al., 2016). For each pathology, the training dataset was sampled such that the ratio of positive and negative cases is $2:1$. The model was trained for 10 epochs with a $41e-$ 4 learning rate. The training process was monitored using the AUC score and the one with the highest validation AUC was kept. Then, the decision threshold $d$ was fine-tuned to optimize a weighted sum of Specificity and Negative Predictive Value (NPV); i.e.,
$d=w_{1}\frac{TN}{TN+FP}+w_{2}\frac{TN}{TN+FN},$ (1)
@@ 4.3 Results
PT3: Referring to Weak Learners
Going beyond our PT2 strategy, we applied our PT3 strategy to further improve diagnostic accuracy. Table 5 provides the TP, FP, and FN prediction counts of LLaVA-Med on the Chexpert test set using the PT2 strategy. Note the large number of FP cases. Hence, we designed our weak learners to suppress FP predictions. Table 6 compares the performance on Chexpert before and after referring to the weak learner. It shows that the F1 prediction accuracy can be substantially increased by introducing weak learner predictions into the prompts. The F1 scores of Cardiomegaly, Edema, and Pleural Effusion increase by 0.115, 0.194 and 0.089, respectively. To further demonstrate the efficacy of our PT3 strategy, Table 7 compares the FP predictions of the PT2 and PT3 strategies. The reduction of FP cases is noteworthy, especially on Edema, for which the FP count is reduced by 78.5% (322).
@@
Table 7: FP counts of LLaVA-Med with the PT2 and PT3 strategies on the Chexpert test set.
@@
Our prompt strategies can also be applied to general domain LVLMs. We studied the performance of LLaVA (Liu et al., 2023b) and MiniGPT-v2 (Zhu et al., 2023) using POPE metrics (Li et al., 2023c), which evaluate the hallucination of LVLMs by asking questions about the presence of objects. The POPE scores of LLaVA and MiniGPT-v2 have high Precision and low Recall. Hence, our weak learner strategy can be used to reduce the FN predictions. We selected an off-the-shelf Fast-RCNN (Girshick, 2015) as the weak learner, fine-tuned the detection threshold of bounding box scores to achieve high Recall, and introduced the positive predictions of the weak learner into the prompts. The results in Table 10 show that the Recall scores across three POPE categories increased by around 7% (Precision scores decrease slightly), thus improving the F1 scores.
@@ 5 Conclusions and Discussion
However, our two strategies are not effective on pathologies with extremely scarce data. For example, providing text explanations for Consolidation, Fracture, Lung Lesion, Pneumonia, and Pneumothorax may not suffice since the visual encoder does not adequately learn meaningful visual features. Moreover, the data may not suffice to adequately train weak learners. A promising direction for future research would be to devise a strategy for handle these rare categories. Retrieval Augmented Generation (RAG) could be a potential solution. In addition to textual explanation of a pathology, for instance, typical example images can be provided to help the model make diagnostic decisions.
A Pathology Explanations

### v3

@@ Abstract
doi:
10.59275/j.melba.2025-1a8b
keywords:
Medical Visual Question Answering, Large Vision-Language Models, Prompt Engineering
volume: 3
@@ 1 Introduction
Research on Large Language Models (LLMs) has yielded astonishing results in recent years. LLMs with billions of parameters have achieved outstanding abilities in a wide range of application scenarios (OpenAI, 2022; OpenAI et al., 2023; Chiang et al., 2023). The success of LLMs has quickly extended into the Vision-Language (VL) domain. Large Vision-Language Models (LVLMs) are built upon LLMs by training adapters that project visual features into tokens that can be interpreted by LLMs (Li et al., 2023b; Zhang et al., 2023a; Liu et al., 2023b). Visual Question Answering (VQA) is an essential skill of LVLMs, and VQA accuracy serves as a test metric for most of these models (Li et al., 2023b; Zhang et al., 2023a; Zhu et al., 2023; Liu et al., 2023b). LVLMs have been pretrained on medical datasets (Li et al., 2023a; Liu et al., 2023c; Singhal et al., 2023) and they have been tested on medical VQA tasks (Lau et al., 2018; He et al., 2020). These Medical LVLMs (MLVLMs) have been able to answer questions regarding the imaging modalities, organs, and abnormalities depicted by the input medical scans.
However, “hallucination” has been a major problem for LVLMs. This refers to the generation of content that is contradictory to the input images. Hallucination can be measured via VQA. One may ask the model questions regarding the existence of objects in the input image(s) and the hallucination level is assessed as the percentage of correctly answered questions. VQA can also potentially serve for medical image diagnosis. Users pose questions regarding a pathology and the MLVLM responds based on its analysis of the medical scans. However, most of the available datasets involve simple questions such as “what is the modality of this image” and “what is the organ/tissue in this image”. MLVLMs have yet to be thoroughly evaluated on VQA accuracy across a broad spectrum of pathologies. Additionally, general VQA models are usually tested by the commonly known accuracy: the percentage of correctly answered questions, which is an unsuitable measure for medical VQA. Medical image classification metrics such as the Precision, Recall, and F1 are more suitable for the evaluation of medical VQA models. Several strategies have been explored to enhance the question answering of LLMs/LVLMs, including chain-of-thought prompting (Zheng et al., 2023), self-consistency (Wang et al., 2023), and retrieval-based augmentation (Caffagni et al., 2024). All these methods involve fine-tuning the models, which is expensive. Training-free methods to improve the VQA accuracy of MLVLMs are desirable.
@@
    We introduce a low-cost weak learner model as a reference for LLaVA-Med, and this effectively reduces the false positive answers.
@@ 2 Related Work
LVLMs are built upon LLMs. A pretrained visual encoder extracts the visual features and an adapter module projects the extracted features to ones that can be understood by the LLM. Models of this type include those by Liu et al. (2023b), Zhu et al. (2023), and Zhang et al. (2023a). During training, the visual encoder and the LLM are usually fixed. VQA is an essential skill of LVLMs. Given an input image, the models should be able to answer questions correctly regarding that image.
@@
Strategies to mitigate hallucination in LVLMs mainly fall into two categories: prompt engineering and model improvement. Prompt engineering is a well developed technique in the natural language domain that provides LLMs with instructions and/or additional information to perform tasks. Vu et al. (2024) and Peng et al. (2023) improve the LLM performance by providing external information. Regarding the former, Liu et al. (2023a) leverage visual instructions constructed from the bounding box information in the input image to prompt LLMs. Zheng et al. (2023) use a chain of thought scheme to prompt the models to perform step-by-step visual-language reasoning like humans, which eventually leads to the correct answers. Wang et al. (2023) generate multiple chains of thought and use the one with the majority vote as the answer. Caffagni et al. (2024) prompt the model with explanations of the terms in questions. Wang et al. (2024) retrieve external knowledge to assist the model in VQA tasks. With regard to the model improvement strategy for reducing hallucination, Sun et al. (2023) improve the visual and text feature alignment through reinforcement learning. Leng et al. (2023) propose a contrastive decoding strategy to reduce reliance on pretrained knowledge. Favero et al. (2024) and Zhao et al. (2024) also focus on the inference stage and propose specialized decoding strategies to mitigate hallucination. Other strategies for reducing hallucination have been proposed. For example, Zhou et al. (2024) design a post-processing model to detect hallucinated objects and rephrase the generated answers, and Sun et al. (2023) adapt a reinforcement learning strategy that uses human evaluation of the hallucination level to improve the model.
@@
There are two approaches to assessing hallucination in LVLMs. The first is VQA. The ground truth information of the input images is leveraged to construct questions regarding the existence of objects in the images (e.g., “Is there a black cat in the image?”), as well as questions about objects which do not exist in the images. The models are evaluated in terms of the percentage of correctly answered questions. Metrics of this type include POPE (Li et al., 2023c), CIEM (Hu et al., 2023), and NOPE (Lovenia et al., 2023). The second approach is to use pre-designed prompts from which the models produce various generations that are then evaluated. Examples include CHAIR (Rohrbach et al., 2018), which counts the hallucinated objects in generated image captions, and MMHAL-BENCH (Sun et al., 2023), which uses GPT-4 (OpenAI et al., 2023) to compare the generations with human answers and determine the propensity toward hallucination.
@@
For MLVLMs, given a medical scan, models such as LLaVA-Med (Li et al., 2023a) and Med-PALM (Liu et al., 2023c) are able to answer questions regarding the types of modalities, the scanned organs, and medical indicators such as opacity. They have demonstrated good performance on medical VQA datasets such as VQA-RAD (Lau et al., 2018), SLAKE (Liu et al., 2021), and Path-VQA (He et al., 2020). XrayGPT (Thawakar et al., 2023) improved model performance by performing an additional round of training on selected high-quality data. CheXagent (Chen et al., 2024) further developed the training process such that the visual encoder, projection layer, and the whole model are trained separately in three steps. However, most medical questions in existing datasets are simple (e.g., view classification). MLVLMs have not yet been tested on a broader range of complex pathologies. Recently, RaDialog (Pellegrini et al., 2023b) trains a separate image classifier and fine-tunes the MLVLM with the classification results on their designed datasets including various medical tasks. The image classifier helps the model generate medical reports with high medical correctness.
@@
Medical Image Classification via VLMs
The most notable vision-language model for medical image classification is ConVIRT (Zhang et al., 2022), which employed a contrastive learning approach to pretrain the model for various tasks. It inspired CLIP (Radford et al., 2021), which is able to utilize the pretrained model for zero-shot classification by searching for the best match between the image features and text features of disease categories. Several CLIP variants, such as BioMedCLIP (Zhang et al., 2023b), ChexZero (Tiu et al., 2022), MedCLIP (Wang et al., 2022), Xplainer (Pellegrini et al., 2023a), Seibold et al. (2022), and Jang et al. (2022), perform well on medical image classification tasks. They can also be fine-tuned for impressive performance on other tasks, such as segmentation and report generation. Compared with LVLMs, they are single-purpose models rather than generative AIs; however, these models may be used as the backbones of encoders in MLVLMs.
@@ 3.1 Prompting With Detailed Explanations
Prompt templates for a number of pathologies are listed in Appendix A.
Figure 3: An example of prompting an MLVLM for medical VQA using both pathology explanations and reference predictions from a weak learner.
@@ 3.2 Prompting With Detailed Explanations and Weak Learners
One can nevertheless enable MLVLMs to benefit by leveraging small models trained on re-sampled datasets. Our method resembles multiagent LLM systems, such as Du et al. (2023), where multiple LLMs debate each other and hallucination can be corrected by referring to the generated outputs of other models. Given that traditional image classifiers are smaller, it is feasible to train multiple small classifiers each of which is trained on re-sampled datasets of a particular pathology. Those models can be further fine-tuned to optimize a single aspect, such as fewer False Positives (FPs) or fewer False Negatives (FNs). The classifiers are applied to the medical images and return preliminary predictions. These predictions are selectively included in the prompts as references for the MLVLM. Hence, MLVLMs can benefit indirectly from the nuanced understanding that these specialized models can provide. This method is meaningful because clinicians usually must balance the trade-off between overtreatment and undertreatment when making healthcare decisions. For instance, they may prefer models having a low FP rate if the cost of overtreatment is higher than that of undertreatment.
@@ 4.1 Datasets
For our main testing regimen, we selected the five pathologies in the Chexpert Competition (Irvin et al., 2019): Atelectasis, Cardiomegaly, Consolidation, Edema, and Pleural Effusion.
@@ 4.2 Implementation Details
Table 3: The Prompt Templates (PTs). {target} is the pathology cited in the questions. {explanation} contains a pathology explanation among those listed in Appendix A. {n} is the probability associated with the weak learner.
@@
As will be justified by our experiments, our weak learner is designed to suppress FP predictions. To this end, we use the pretrained ResNet50 (He et al., 2016). Given the low cost of the weak learner, we train a model separately for each pathology with each training dataset sampled such that the ratio of positive and negative cases is $2:1$. The model was trained for 10 epochs with a $41e-$ 4 learning rate. The training process was monitored using the AUC score and the one with the highest validation AUC was kept. Then, the decision threshold $d$ was fine-tuned to optimize a weighted sum of Specificity and Negative Predictive Value (NPV); i.e.,
$d=w_{1}\frac{\text{TN}}{\text{TN}+\text{FP}}+w_{2}\frac{\text{TN}}{\text{TN}+\text{FN}},$ (1)
@@ 4.3 Results
Pathology AUC Precision Recall F1
Atelectasis 82.2 62.0 56.7 59.2
Cardiomegaly 85.0 74.4 38.3 50.6
Consolidation 81.7 100 2.9 5.6
Edema 87.3 46.5 61.2 54.2
Pleural Effusion 91.2 36.6 94.2 52.7
Table 6: Performance of the weak learner on the Chexpert test sets for the 5 pathologies.
PT3: Referring to Weak Learners
Going beyond our PT2 strategy, we applied our PT3 strategy to further improve diagnostic accuracy. Table 5 provides the TP, FP, and FN prediction counts of LLaVA-Med on the Chexpert test set using the PT2 strategy. Note the large number of FP cases. Hence, we designed our weak learners to suppress FP predictions. As mentioned in Section 4.2, we trained the model for the highest AUC and then fine-tuned the decision threshold. Table 6 reports the AUC, Precision, Recall and F1 (after fine-tuning) of the weak learner. It is important to note that we fine-tuned the decision threshold to achieve high specificity and negative predictive value; thus, the reported Precision, Recall, and F1 scores are based on this fine-tuned threshold and may not be directly comparable to those of other models. Table 7 compares the performance on Chexpert before and after referring to the weak learner. It shows that the F1 prediction accuracy can be substantially increased by introducing weak learner predictions into the prompts. The F1 scores of Cardiomegaly, Edema, and Pleural Effusion increase by 0.115, 0.194 and 0.089, respectively. To further demonstrate the efficacy of our PT3 strategy, Table 8 compares the FP predictions of the PT2 and PT3 strategies. The reduction of FP cases is noteworthy, especially on Edema, for which the FP count is reduced by 78.5% (322).
@@
Additional VQA Experiments
SOTA Benchmark
@@
Table 8: False positive counts of LLaVA-Med with the PT2 and PT3 strategies on the Chexpert test set.
@@
Our prompt strategies can also be applied to general domain LVLMs. We studied the performance of LLaVA (Liu et al., 2023b) and MiniGPT-v2 (Zhu et al., 2023) using POPE metrics (Li et al., 2023c). POPE evaluates the hallucination of LVLMs by asking questions about the existing/non-existing objects on given images. The input images are from MSCOCO dataset and there are three questions categories: Random (ramdom sample objects for questions), Popular (frequent objects), and Adversarial (frequent but non-existent objects). The performance is evaluated by the Precision, Recall and F1 of correctly answered questions. The POPE scores of LLaVA and MiniGPT-v2 have high Precision and low Recall. Hence, our weak learner strategy can be used to reduce the FN predictions. We selected an off-the-shelf Fast-RCNN (Girshick, 2015) as the weak learner, fine-tuned the detection threshold of bounding box scores to achieve high Recall, and introduced the positive predictions of the weak learner into the prompts. The results in Table 11 show that the Recall scores across three POPE categories increased by around 7% (Precision scores decrease slightly), thus improving the F1 scores.
@@ 5 Conclusions and Discussion
However, our two strategies are not effective on pathologies with extremely scarce data. For example, providing text explanations for Consolidation, Fracture, Lung Lesion, Pneumonia, and Pneumothorax may not suffice since the visual encoder does not adequately learn meaningful visual features. Moreover, the data may not suffice to adequately train weak learners. A promising direction for future re-
search would be to devise a strategy for handling these rare categories. Retrieval Augmented Generation (RAG) could be a potential solution. For instance, in addition to textual explanations of pathologies, typical example images can be provided to help the model make diagnostic decisions.
[…]
DG was supported in part by an unrestricted donation to UCLA from VoxelCloud, Inc.
[…]
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
[…]
We have no conflicts of interest.
Data Availability
Only publicly available datasets were used.
Appendix A Pathology Explanations